# SMooGPT: Stylized Motion Generation using Large Language Models

### v1

@@ Appendix D Design Details of Competitors
Since no existing framework directly addresses the problem of text-guided motion stylization, we develop baseline approaches that integrate ChatGPT with existing motion generation methods, MotionGPT (Jiang et al., 2023) and MLD (Chen et al., 2023). Specifically, we follow the same procedure used in dataset construction to compose the content text and style label into a unified global description, which is then fed into the two motion generation models.
@@
We leverage the source code of SMooDi to retrian in the in-domain subset. Because there is a trade-off between content preservation and style reflection when using two types of style guidance, we set the classifier-free guidance weight to 1 and retain the original weight for classifier-based guidance during inference to better preserve content, aligning more closely with the original paper.
Table A1. Quantitative results of pretraining and instruction tuning on the HumanML3D dataset.
Method FID $\downarrow$ MM Dist $\downarrow$ R-Pre.(Top-3) $\uparrow$ FS-Ratio $\downarrow$
SMooGPT(Pre-trained) 0.433 3.124 0.754 0.0982
SMooGPT(Fine-tuned) 0.396 3.125 0.761 0.0924
Appendix E Additional Results
The results of pretraining and instruction
We report the performance of pretraining and instruction tuning on HumanML3D in Table A1.

### v2

@@ Appendix E Design Details of Competitors
Since no existing framework directly addresses the problem of text-guided motion stylization, we develop baseline approaches that integrate ChatGPT with existing motion generation methods, MotionGPT (Jiang et al., 2023), MLD (Chen et al., 2023) and ParCo (Zou et al., 2024). Specifically, we follow the same procedure used in dataset construction to compose the content text and style label into a unified global description, which is then fed into the three motion generation models.
@@
We leverage the source code of SMooDi to retrian in the two settings. For the zero-shot setting, we early-stop training at $10$ epochs to prevent content forgetting caused by over-focusing on the reference motion. For the held-out setting, we follow the original SMooDi training protocol and train the model for $50$ epochs. To balance the trade-off between content preservation and style reflection, we set the classifier-free guidance weight to $1$. Furthermore, we disable classifier-based guidance during inference, as it relies on a pre-trained encoder that fails to generalize to unseen reference motions.
Figure A1. Failure Cases. Representative failure cases illustrating the limitations of our method. Left: For motions requiring sharp temporal transitions (e.g., a sharp salute), the model captures the overall arm trajectory but fails to produce an abrupt stopping phase. Right: For out-of-distribution motion patterns (e.g., a bird-flying walk), the generated motion remains constrained to the learned motion manifold, resulting in missing periodic wing-flapping dynamics.
Appendix F User Evaluation Details
We conduct a user perceptual study using pairwise comparisons to evaluate text-guided motion stylization. In each test case, participants are presented with two motion clips generated by our method and a competing approach, respectively, given the same motion content text and motion style text. To ensure a fair comparison, all videos are anonymized and randomized, and no information about the generating method is revealed to the participants.
Participants are asked to evaluate each motion pair based on three criteria: content preservation, style reflection, and motion realism. For content preservation, participants assess how well the generated motion matches the semantics described in the motion content text. For style reflection, they evaluate how clearly the target motion style is expressed in the generated animation. For motion realism, participants judge the naturalness and physical plausibility of the motion, focusing on movement continuity and visual coherence.
A total of 21 motion pairs are evaluated, evenly divided into three groups corresponding to comparisons with three competing methods. We invite 75 participants to complete the study, which is conducted online using Google Forms. Each motion pair is presented side by side, and participants are required to indicate their preference for each of the three criteria independently.
Appendix G Limitations and Future Work
While our method demonstrates strong performance, it remains subject to limitations imposed by the intermediate representation, as illustrated by the representative failure cases in Fig. A1. First, the discrete nature of the body-part text space lacks the temporal granularity required to fully capture sharp and transient motion dynamics. As shown in Fig. A1(a), although the model correctly generates the overall arm trajectory for a sharp salute, it fails to produce an abrupt stopping phase, resulting in an overly smooth transition near the motion peak. This limitation arises from the fact that discrete, text-based descriptions cannot precisely specify fine-grained temporal phases such as instantaneous stops or rapid impulse changes. Second, while the LLM exhibits strong reasoning capabilities for generalizing to novel styles, the final generation quality is fundamentally bounded by the representational capacity of the body-part VQ-VAE. As illustrated in Fig. A1(b), for out-of-distribution motion patterns such as a bird-flying walk, the model captures the intended spatial configuration (e.g., extended arms) but fails to synthesize clear and periodic wing-flapping dynamics. This behavior reflects the constraint that the decoder can only generate motions within the learned motion manifold, limiting its ability to extrapolate to motion patterns that are underrepresented or absent in the training data.
In future work, we aim to address these constraints by enriching the intermediate representation. Specifically, we plan to augment body-part textual descriptions with a series of keypose-based tokens that explicitly encode spatial configurations over time. Such temporally-aware extensions will enable the model to capture motion dynamics more effectively, offering a richer and more precise representation for both content and style.
Appendix H Additional Results
@@
Example: Global-to-Body-Part Decomposition.
We provide a concrete example illustrating how the model decomposes a global motion description into fine-grained body-part-centric texts following the instruction template above.
Input (Global Motion Text):
"A person bends their left arm at the elbow and raises their right arm towards it, then lowers both arms."
Output (Body-Part Texts):
Root: The body remains mostly upright with a slight weight shift to maintain balance.
Backbone: The backbone stays straight and aligned, providing support for the arm movements.
Left Arm: The left arm bends at the elbow and serves as the focal point of the motion.
Right Arm: The right arm raises upward toward the left arm with controlled movement, then lowers.
Left Leg: The left leg provides stable support throughout the motion.
Right Leg: The right leg maintains a steady position with minimal movement.
Example: Content–Style Composition.
We provide an example illustrating how the model composes motion content and motion style descriptions in the body-part space to produce a unified and coherent representation.
Input (Motion Content Body-Part Texts):
Root: Moves forward with a steady walking motion.
Backbone: Remains upright to support forward locomotion.
Left Arm: Moves forward naturally as part of walking.
Right Arm: Moves forward naturally as part of walking.
Left Leg: Steps forward to initiate locomotion.
Right Leg: Steps forward to continue the walking motion.
Input (Motion Style Body-Part Texts: InTheDark):
Root: Moves cautiously with a restrained and steady gait.
Backbone: Maintains a controlled and composed posture.
Left Arm: Extends forward at chest height with minimal swing.
Right Arm: Extends forward at chest height with minimal swing.
Left Leg: Takes deliberate, measured steps.
Right Leg: Takes deliberate, measured steps.
Output (Composed Body-Part Texts):
Root: Moves forward with a steady and cautious gait, maintaining balance and coordination.
Backbone: Maintains stability and alignment throughout the motion.
Left Arm: Extends straight ahead at chest height, held steady while walking.
Right Arm: Extends straight ahead at chest height, held steady while walking.
Left Leg: Steps forward with a large diagonal movement to the front right.
Right Leg: Steps forward with a large diagonal movement to the front right.